If the voltage between the emitter and base is 0.6V and the transmitter temperature is 290.5K, then calculate Ic/Io ratio.
Answer #1

Ic = Io(evbe/vt-1)

Vt = kt/q

K is bolzman constant = 1.3806488*10-23J/K

t is temperature =290.5K

q is charge of the electron =1.602*10-19C

Vt = 0.0211

Ic/Io = e0.6/00.0211-1 = 2.23*1012